OpenStudy (anonymous):

find the indefinite integral of (sinx)^3 * cos(x)

OpenStudy (anonymous):

\[\int\limits_{}^{}u ^{3}du=?\]

OpenStudy (anonymous):

integral of (sinx)^3 * cos(x)dx let u=sin x du=cosx

OpenStudy (anonymous):

\[\int\limits_{}^{}u ^{3}du=\frac{ 1 }{ 4 } u ^{4}+C\]

OpenStudy (anonymous):

=\[\frac{ 1 }{ 4 }(\sin x)^{4}+C\]

OpenStudy (anonymous):

yeah, thnks i was being retarded